# A is for Absorption: Studying Feature Splitting and Absorption in Sparse Autoencoders

### v1

@@ 1 Introduction
The ability to decompose LLM activations into monosemantic latents presents an opportunity to not only interpret but also modify internal model representations for various downstream tasks (Marks et al., 2024). However, despite these theoretical advantages, most demonstrations of SAEs for interpretability have been limited to toy models (Elhage et al., 2022) or concepts that are difficult to validate empirically (Huben et al., 2024), where precision and recall are hard to verify.
@@ 3 Experimental setup
To determine the causal effect of SAE latents on the first-letter identification task we conduct ablation studies. We use a metric consisting of the logit of the correct letter token minus the mean logit of all incorrect letters. This measures the propensity of the model to choose the correct starting letter as opposed to other letters. Formally, our metric $m$ is defined below, where $g$ refers to the final token model logits, $L$ is the set of uppercase letters, and $y$ is the uppercase letter that is the correct starting letter:
$m=g[y]-\frac{1}{|L|-1}\sum_{l\in\{L\setminus y\}}g[l]$
@@ 5.3 Decomposing SAE Latents
The phenomenon of SAEs of different sides splitting a feature into various smaller latents was first described in Bricken et al. (2023), which noted that different SAE widths and sparsities induce latents of different granularity, with wider SAEs often learning more specific variants of features in narrower SAEs.
@@ 6 Discussion
The significance of our results hinges in part on whether we believe that using SAE latent classification performance on a simple first-letter identification task is a valid measure of their interpretability. One may argue it is unreasonable to judge an SAE latent against a linear probe directly optimized to perform the same classification task and that it has been established that sometimes unsupervised methods can surprise us (Nanda, 2023). However, we argue that for a latent to be considered interpretable, its behavior should match what one would reasonably expect the latent to be doing after inspecting its activation patterns. In our experiments, we use SAE latents that do appear to perform first letter classification, and then evaluate how well they perform this task. We validate as well that these latents causally mediated model performance on the first-letter task in Appendix A.2. We have been convinced that these latents should reasonably be considered “first-letter” latents and thus that their performance on the first-letter identification task is a valid measure of their interpretability.
@@
In trying to understand why SAE latents fail to match the performance of LR probes, we identified a form of feature splitting we call “feature absorption”. Feature absorption may be particularly problematic for SAEs because it creates an interpretability illusion where we believe we have found an interpretable latent, but absorption induces lower recall by creating clear false negatives / exceptions to the mainline interpretation of the latent. This lower recall poses problems for methods which rely on using SAEs to find sparse circuits (Marks et al., 2024), as the number of latents needed to characterize model behavior may be much larger than expected. For example, we may believe we have found a SAE latent which tracks deceptive behavior in the model, but due to feature absorption, there may be many cases where that latent fails to fire with an absorbing feature firing instead. We find that feature absorption happens even in high-recall latents, so this is not only a problem for low L0 SAEs and appears to be a more fundamental issue.
Feature absorption is not predicted by toy-models where features do not have high co-occurrence with each other (Elhage et al., 2022). While more work is required to understand why feature absorption occurs, we suspect it is a consequence of co-occurrence between sparse and dense features. If a dense feature like “starts with letter D” always co-occurs with a more sparse feature like “dogs”, the SAE can increase sparsity by absorbing the “starts with D” feature into a “dogs” latent.
It remains to be seen if we can predict or identify excepted instances where a feature “should have activated” but does not activate due to absorption. One promising direction is meta-SAEs, a novel method for decomposing SAE latents and may decompose absorbed features (Bussmann et al., 2024). One way to interpret our results is that competition may exist between “latents” and “meta-latents” for activation on particular examples and that re-allocation of examples between SAE latents enables SAEs to interpolate between different possible decompositions.
@@
We hope as well this investigation may lead to research into solutions, particularly those involving Meta-SAEs (Bussmann et al., 2024), to solve or mitigate feature absorption. One solution may be attribution dictionary learning (Olah et al., 2024).
@@ Appendix A Appendix
A.1 How good is Gemma-2 on character identification tasks?
@@
A.2 Intervening on the first letter
@@
A.3 Probe cosine similarity vs k=1 sparse probing
@@
A.4 Precision, recall, and F1 score for the first-letter task
@@
A.5 Causal interventions and absorption

### v2

@@ 1 Introduction
The ability to decompose LLM activations into monosemantic latents presents an opportunity to not only interpret but also modify internal model representations for various downstream tasks (Marks et al., 2024). However, despite these theoretical advantages, most demonstrations of SAEs for interpretability have been limited to toy models (Sharkey et al., 2022) or concepts that are difficult to validate empirically (Huben et al., 2024). Bricken et al. (2023) performs some case-studies on SAE latents in a one-layer transformer model, but assumes that gaps in recall are minor and due to imperfections in the model or data and does not investigate further.
@@ 3 Experimental setup
To determine the causal effect of SAE latents on the first-letter identification task we conduct ablation studies. We use a metric consisting of the logit of the correct letter minus the mean logit of all incorrect letters. This measures the propensity of the model to choose the correct starting letter as opposed to other letters. Formally, our metric $m$ is defined below, where $g$ refers to the final token logits, $L$ is the set of uppercase letters, and $y$ is the uppercase letter that is the correct starting letter:
$m=g[y]-\frac{1}{|L|-1}\sum_{l\in\{L\setminus y\}}g[l]$
@@ 5.3 Decomposing SAE Latents
The phenomenon of SAEs of different sides splitting a feature into various smaller latents was first described in Bricken et al. (2023), which noted that different SAE widths and sparsities induce latents of different granularity, with wider SAEs often learning more specific variants of features.
@@ 6 Discussion
One may argue it is unreasonable to judge an SAE latent against a linear probe directly optimized to perform the same classification task and that it has been established that sometimes unsupervised methods can surprise us (Nanda, 2023). However, we argue that for a latent to be considered interpretable, its behavior should match what one would reasonably expect the latent to be doing after inspecting its activation patterns. In our experiments, we use SAE latents that do appear to perform first letter classification, and then evaluate how well they perform this task. We validate as well that these latents causally mediated model performance on the first-letter task in Appendix A.3. We are convinced that these latents should reasonably be considered “first-letter” latents and that their performance on the first-letter identification task is a valid measure of their interpretability.
@@
In trying to understand why SAE latents fail to match the performance of LR probes, we identified a form of feature splitting we call “feature absorption”. Feature absorption may be particularly problematic for SAEs because it creates an interpretability illusion where we believe we have found an interpretable latent, but absorption induces lower recall by creating clear false negatives to the mainline interpretation of the latent. For example, we may believe we have found a SAE latent which tracks deceptive behavior in the model, but due to feature absorption, there may be many cases where that latent fails to fire. This lower recall poses problems for methods which rely on using SAEs to find sparse circuits (Marks et al., 2024), as the number of latents needed to characterize model behavior may be much larger than expected. We find that feature absorption happens even in high-recall latents, so this is not only a problem for low L0 SAEs and appears to be a more fundamental issue.
Feature absorption is not predicted by toy-models where features do not co-occur (Bricken et al., 2023). While more work is required to understand why feature absorption occurs, we suspect it is a consequence of co-occurrence between sparse and dense features. If a dense feature like “starts with letter D” always co-occurs with a more sparse feature like “dogs”, the SAE can increase sparsity by absorbing the “starts with D” feature into a “dogs” latent.
It remains to be seen if we can predict or identify excepted instances where a feature “should have activated” but does not activate due to absorption. One promising direction is meta-SAEs, a novel method for decomposing SAE latents and may decompose absorbed features (Bussmann et al., 2024). One interpretation of our results is that competition may exist between “latents” and “meta-latents” for activation on particular examples and that re-allocation of examples between SAE latents enables SAEs to interpolate between different possible decompositions.
@@
We hope as well this investigation may lead to research into solutions, particularly those involving Meta-SAEs (Bussmann et al., 2024), to solve or mitigate feature absorption. Anoter possible solution may be attribution dictionary learning (Olah et al., 2024).
@@ Appendix A Appendix
A.1 Glossary of Terms
Sparse Autoencoders (SAEs):
Neural networks trained to reconstruct their input while enforcing sparsity in their hidden layer. In the context of this paper, SAEs are used to decompose the dense activations of language models into more interpretable features.
SAE error term:
When inserting a SAE into the computation path of the model, errors in SAE reconstruction will propagate to later parts of the model and can change the model output. We refer to the error as the SAE error term, and corresponds to the difference between the SAE output and the original SAE input activation. Marks et al. (2024) introduced the idea of adding this error term back to the SAE output to ensure that the SAE does not change model output.
Latent:
We refer to neurons in the hidden layer of a SAE as latents to avoid overloading the term “feature”. This is in contrast to earlier work which used the term “feature” to refer to both human-interpretable concepts and SAE hidden layer neurons.
Feature:
We use the term “feature” to refer to an idealized human-interpretable concept that the model represent in its activations and which a SAE latent may or may not represent.
Monosemantic:
Referring to a feature or representation that corresponds to a single, clear semantic concept. In the context of SAEs, a monosemantic feature would ideally capture one interpretable aspect of the input.
Interpretable:
A latent being interpretable is not well defined in the field, making it difficult to ensure that different authors mean the same thing when referring to SAE interpretability. When we refer to a SAE latent as being interpretable in this work, we mean that it should behave in line with how it appears to behave after inspecting its activation patterns. If an SAE latent appears to track a feature X by a reasonable inspection of its activations but has subtle deviations from this behavior in reality, we say this is not interpretable. We thus measure interpretability via classification performance when a latent appears to be a classifier over some feature.
Feature dashboard:
A dashboard showing activation patterns and max-activating examples for a SAE latent. Feature dashboards are commonly used to interpret the behavior of an SAE latent.
Neuronpedia:
A platform, https://neuronpedia.org, which hosts feature dashboards for popular SAEs (Lin & Bloom, 2023).
Feature splitting:
A phenomenon in SAEs introduced by Bricken et al. (2023), where a SAE latent tracks a general feature in a narrow SAE, but splits into multiple more specific SAE latents in a wider SAE. For instance, a latent tracking “starts with L” in a narrow SAE may split into a latent tracking “starts with capital L” and a latent tracking “starts with lowercase L” in a wider SAE.
Feature absorption:
A problematic form of feature splitting where a SAE latent appears to track an interpretable feature, but that latent has seemingly arbitrary exception cases where it fails to fire. Instead, an approximately token-aligned feature “absorbs” the feature direction and fires in place of the main latent.
Linear probe:
A simple linear classifier (typically logistic regression) trained on the hidden activations of a neural network to predict some property or task. Used to assess what information is linearly decodable from the network’s representations.
K-sparse probing:
A variant of linear probing where only the k most important features (as determined by some selection method) are used to train the probe. This helps identify which specific neurons or features are most relevant for a given task.
Ablation study:
An experimental method where a component of a system (in this case, a neuron or feature in a neural network) is removed or altered to observe its effect on the system’s performance. This helps determine the causal importance of the component.
Integrated gradients (IG):
An attribution method that assigns importance scores to input features by accumulating gradients along a path from a baseline input to the actual input. In this paper, it’s used as an approximation technique for ablation studies.
In-context learning (ICL):
A paradigm where a language model is given examples of a task within its input prompt, allowing it to adapt to new tasks without fine-tuning. Often used with few-shot learning techniques.
Residual stream:
In the context of transformer architectures, the residual stream refers to the main information flow that bypasses the self-attention and feed-forward layers through residual connections.
Logits:
The raw, unnormalized outputs of a neural network’s final layer, before any activation function (like softmax) is applied. In language models, logits typically represent the model’s scores for each token in the vocabulary.
Activation patching:
An interpretability technique where activations at specific locations in a neural network are replaced or modified to observe the effect on the network’s output. This helps in understanding the causal role of different parts of the network in producing its final output.
A.2 How good is Gemma-2 on character identification tasks?
@@
A.3 Intervening on the first letter
@@
A.4 Probe cosine similarity vs k=1 sparse probing
@@
A.5 Precision, recall, and F1 score for the first-letter task
@@
A.6 Causal interventions and absorption